The dimensions of a rectangle are 6cm by 8cm. When both dimensions are increased by equal amounts, the area of the rectangle is increased by 208 square cm. Find the dimensions of the new rectangle.
Answer by checkley75(3666)

(6+X)(8+X)=6*8+208
48+14X+X^2=48+208
X^2+14X-208=0
USING THE QUADRATIC EQUATION WE GET:
X=(-14+-SQRT[14^2-4*1*-208])/2*1
X=(-14+-SQRT[196+832])/2
X=(-14+-SQRT10280/2
X=(-14+-32.06)/2
X=(-14+32.06)/2
X=18.06/2
X=9.03 ANSWER.
PROOF
(6+9.03)(8+9.03)=48+208
15.03*17.03=256
256=256